Boston is 295 miles from new York along a certain route. a car starts from boston at 1 pm and travels along this route toward new York at a steady rate of 50 miles per hour.
Another car starts from new York at 1:30 pm and travels along this route toward boston at a steady rate of 40 miles per hour. at what time do the cars pass each other?

Found 2 solutions by ankor@dixie-net.com, greenestamps:
Answer by ankor@dixie-net.com(22740)

Boston is 295 miles from new York along a certain route. a car starts from boston at 1 pm and travels along this route toward new York at a steady rate of 50 miles per hour.
Another car starts from new York at 1:30 pm and travels along this route toward boston at a steady rate of 40 miles per hour. at what time do the cars pass each other?
:
Change 30 min to .5 hrs
let t = travel time of the 1st car when they meet
let (t-.5) = travel time of the 2nd car
:
When they meet, the sum of their distances will be 295 mi
Write dist equation; dist = speed * time
50t + 40(t-.5) = 295
50t + 40t - 20 = 295
90t = 295 + 20
90t = 315
t = 315/90
t = 3.5 hrs is the travel time of the 1st car
therefore
1:00 + 3.30 = 4:30 is the time they meet
:
:
Check this; find the actual dist each traveled, car 2 traveled for 3 hrs
3.5(50) = 175 mi
3(40) = 120 mi
------------------
total dist: 295 mi


Answer by greenestamps(13198)



The first car travels for half an hour (30 minutes) at 50mph before the second car starts. That's 25 miles; 270 miles are left.

After the second car starts, the two cars are moving towards each other at a combined 90mph. 270 miles at 90mph takes 270/90 = 3 hours.

So the cars pass each other 3 hours after the second car starts -- at 4:30pm.